### v1

# UStyle: Waterbody Style Transfer of Underwater Scenes by Depth-Guided Feature Synthesis

Abstract
The concept of waterbody style transfer remains largely unexplored in the underwater imaging and vision literature. Traditional image style transfer (STx) methods primarily focus on artistic and photorealistic blending, often failing to preserve object and scene geometry in images captured in high-scattering mediums such as underwater. The wavelength-dependent nonlinear attenuation and depth-dependent backscattering artifacts further complicate learning underwater image STx from unpaired data. This paper introduces UStyle, the first data-driven learning framework for transferring waterbody styles across underwater images without requiring prior reference images or scene information. We propose a novel depth-aware whitening and coloring transform (DA-WCT) mechanism that integrates physics-based waterbody synthesis to ensure perceptually consistent stylization while preserving scene structure. To enhance style transfer quality, we incorporate carefully designed loss functions that guide UStyle to maintain colorfulness, lightness, structural integrity, and frequency-domain characteristics, as well as high-level content in VGG and CLIP (contrastive language-image pretraining) feature spaces. By addressing domain-specific challenges, UStyle provides a robust framework for no-reference underwater image STx, surpassing state-of-the-art (SOTA) methods that rely solely on end-to-end reconstruction loss. Furthermore, we introduce the UF7D dataset, a curated collection of high-resolution underwater images spanning seven distinct waterbody styles, establishing a benchmark to support future research in underwater image STx. The UStyle inference pipeline and UF7D dataset is released at: https://github.com/uf-robopi/UStyle.
Index Terms:
Style Transfer; Waterbody Fusion; Underwater Imaging; Underwater Vision.
I Introduction
Image style transfer (STx) in underwater imagery has significant potential in data augmentation, robotic vision, photometry, and imaging technologies [1]. Traditional neural style transfer (NSTx) approaches [2, 3] primarily focus on terrestrial imagery, aiming to transfer artistic or photometric features across images. However, the concept of “waterbody style” transfer presents a distinct challenge [1], where the goal is to generate an underwater image in the style of a different aquatic environment while preserving the structural integrity of the scene. Although unpaired images from various waterbodies can be collected, conventional NSTx techniques—optimized for perceptual photorealism [4] or artistic blending [5, 6, 7]—struggle to accommodate the unique distortions introduced by underwater light absorption and scattering [8]. As a result, these methods often produce unnatural style adaptation and significant structural loss.
Currently, no learning-based frameworks exist for waterbody STx in underwater imagery. Our previous work, AquaFuse [1], introduced a physics-based approach for waterbody fusion, facilitating data augmentation across different aquatic environments. However, this method relies on a closed-form formulation that requires a carefully selected, noise-free reference image and prior scene information, including depth, incident angle, and attenuation parameters. Consequently, leveraging large-scale vision-based learning models for robust underwater STx without explicit reference images or prior information remains a challenge. Recent advancements in NST [4, 7] and vision-language models [9, 10] present an opportunity to address this gap. However, developing a learning-driven approach capable of accurately adapting underwater distortions while preserving structural details remains an unexplored area of research.
In this paper, we propose UStyle, the first data-driven learning pipeline for waterbody STx in underwater imagery. We present the model architecture, choice of loss functions, and the training strategy required to address the limitations of existing NSTx pipelines for underwater STx. Specifically, UStyle integrates a residual encoder-decoder to extract local structural details at different spatial resolutions while leveraging a progressive multi-stage training strategy to refine waterbody feature learning by domain-specific loss functions. Moreover, to guide the stylization process, we use a physics-based method [1] to synthesize the waterbody from the style image by following the state-of-the-art (SOTA) underwater image formation model [8]. This synthesized waterbody style is then fused by a novel depth-guided whitening and coloring transform (DA-WCT) mechanism, which makes sure that no object/scene information is transferred to the content image. A few sample examples are illustrated in Fig. 1.
Figure 1: Visualizations of the underwater image STx capabilities of UStyle are shown. The content image (top left) is transformed into different waterbody styles. For any given style image, note that its waterbody characteristics are transferred to the content scene, while its object geometry and structural details of the content image remain preserved. Such no-reference waterbody fusion is useful as a geometry-preserving data augmentation tool in underwater imaging, photometry, and robotics applications.
For comprehensive evaluation, we compile an underwater dataset named UF7D – featuring $4,050$ high-resolution images with paired depth annotations across seven different waterbody styles: Blue (B), Green (G), Deep Blue (DB), Deep Green (DG), Clear (C), Blue-Green (B-G), and Green-Blue (G-B). It enables data-driven learning for underwater STx, allowing the model to generalize across diverse aquatic environments with varying optical properties. Most images are sourced from existing SOTA databases [11, 12, 13, 14, 15] for imaging, object detection, image enhancement, and robotics applications; thus, it serves as a useful benchmark for future research. To this end, we conducted a thorough benchmark evaluation of underwater STx, by comparing UStyle with several SOTA methods named StyTR-2 [7], MCCNet [16], ASTMAN [17], IEContraAST [5], PhotoWCT2 [4], and AquaFuse [1].
The key contributions of this work are summarized below:
  1. $1$)
    Underwater STx learning pipeline and dataset: UStyle is the first data-driven STx framework for underwater scenes, eliminating the need for explicit scene priors required by physics-based STx models [1]. Instead of relying on carefully selected reference images and waterbody parameters, UStyle employs a learning-based STx approach that robustly adapts to diverse aquatic environments with varying waterbodies. We also introduce the UF7D dataset, a comprehensive collection of $4,050$ high-resolution underwater images of waterbody styles, which supports supervised training and benchmark evaluation across varying underwater conditions.
  2. $2$)
    Novel contributions in UStyle: We design and develop UStyle to (learn to) blend image features while being aware of the object geometry in underwater scenes for waterbody STx. Our proposed architecture (a) integrates a ResNet-based [18] encoder-decoder with hierarchical skip connections [19] for effective multi-scale feature fusion; (b) employs a progressive multi-stage training strategy [20] for stabilized style-content blending; and (c) domain-specific loss functions learn image statistics (color, lightness) [21], structural details (SSIM [22], FFT [23]) and high-level image contents (VGG [24], CLIP [25]). This guided training strategy is leveraged by our proposed DA-WCT mechanism for depth-aware waterbody stylization, ensuring perceptual fidelity and geometric consistency.
  3. $3$)
    SOTA performance and benchmark evaluation: We conduct comprehensive experimental evaluations validating that UStyle consistently outperforms leading STx methods—such as StyTR-2 [7], MCCNet [16], ASTMAN [17], IEContraAST [5], and PhotoWCT2 [4]—across key metrics such as PSNR, SSIM, RMSE, and LPIPS. Ablation studies also reveal that our proposed DA-WCT mechanism and domain-specific loss functions can deal with the underwater image distortions to generate perceptually coherent stylization. We also compare the performance with physics-based approaches—such as AquaFuse [1]—to validate that UStyle-generated images preserve the structural contents of the scene while fusing waterbody styles. These results establish UStyle as a robust and effective solution for no-reference underwater image STx applications.
UStyle provides a robust and efficient solution for no-reference underwater STx, addressing the limitations of traditional artistic and photorealistic STx approaches. Its ability to preserve geometric consistency while achieving realistic waterbody transformations makes it an ideal framework for real-world marine robotics and imaging applications. UStyle opens up significant opportunities for further research in real-time video-based stylization, data augmentation, and interactive augmented and virtual reality (AR/VR) domains.
II Background and Related Work
Image style transfer (STx) techniques blend the content of one image with the style of another [26, 27, 28], unlocking a wide array of applications across computer vision, virtual reality, robotics, and digital art creation. We can group current advancements in STx into critical domains such as artistic, photorealistic, non-photorealistic, semantic, and multimodal techniques. Each category uses its methods to improve the adaptability and realism of STx, such as optimization algorithms, feed-forward architectures, attention mechanisms, and transformers (see Fig. 2). These approaches collectively push the boundaries of what is achievable in image synthesis, making it possible to preserve structure, adapt to high-resolution outputs, and achieve detailed, context-aware stylization.
  SOTA Approaches Prominent Literature
  Optimization models [26, 29, 27, 30, 31]
Diffusion-based models [32, 33, 34, 35]
Feed-Forward models [36, 37, 38, 4, 39]
Attention models [40, 41, 42, 43]
Transformers [44, 7, 45, 46, 47, 48, 49]
CNNs and GANs [50, 51, 52, 53, 54]
Physics-based models [55, 56, 57, 58, 59]
Cartoonization models [60, 61, 62, 63, 64]
Texture/pattern analysis [65, 3, 66, 67]
Segmentation models [68, 69, 70, 71, 72, 73]
Region-based models [74, 75, 76, 77]
Multi-style fusion [70, 78, 79, 80]
Few-shot & Zero-shot [81, 82, 83]
Vision-language models [9, 10, 84]
Figure 2: A snapshot of the SOTA literature on image-based style transfer (STx) approaches.
II-A Artistic and Non-Photorealistic Style Transfer
Artistic STx aims to apply one image’s artistic characteristics, such as brushstrokes, textures, or color palettes, to another image’s content. Following the seminal work by Gatys et al. [85], many deep visual learning-based networks have been proposed [36, 37, 38]. In particular, Johnson et al. [86] and Ulyanov et al. [87] develop feed-forward networks that apply predefined styles in real-time, making the process much faster. One limitation of their approach is that their network needs to be trained for a specific style. Huang and Belongie [36] address this issue with adaptive instance normalization (AdaIN). AdaIN adjusts the mean and variance of content features to match style features, enabling the transfer of arbitrary styles [88, 89, 90]. Li et al. [37] introduce the whitening and coloring transform (WCT), which applies a whitening operation to remove content, followed by a coloring operation to inject style statistics.
Non-photorealistic rendering (NPR) methods are a superset of artistic STx, focusing on cartooning, sketches, and more abstract art. Early work in this domain is grounded in traditional computer graphics with handcrafted filters and procedural algorithms, whereas recent work uses deep learning techniques to achieve more sophisticated stylizations. Zhao et al. [60] introduces a GAN-based approach for high-quality image cartoonization, capturing stylistic nuances with remarkable fidelity. Ahn et al. [61] develop an interactive framework that allows users to influence the stylization in real-time, enhancing creativity and personalization. Chen et al. [62] proposes DecoGAN, enabling controllable decoration and stylization of images with user inputs. Additionally, techniques such as language-guided STx [65], fast neural doodling [3], and style consistency models [66] expand the possibilities of NPR, making sophisticated stylization accessible to a broader audience. Kim et al. [67] focuses on texture and pattern analysis to preserve fine-grained details during STx, resulting in more cohesive and visually appealing images. These advancements collectively push the boundaries of NPR, enabling the creation of diverse and artistically rich images through deep learning techniques.
Various attention mechanisms have enhanced artistic STx in recent years by capturing complex relationships between content and style. For instance, self-attention GANs [40] enables different parts of an image to interact more effectively to enhance spatial awareness in the artistic rendering process. Building on this, style-attention networks (SANet) [41] achieve more precise style adjustments in different image areas for refined artistic control. These advancements help balance global and local features, improving the quality of fine-grained stylization and overall composition.
II-B Photorealistic Style Transfer
Photorealistic STx applies stylistic changes while keeping the realism and structure of the original image intact. This is essential in understanding visual scenes and robotics applications. Early methods like deep photo style transfer (DPST) [29] use optimization techniques to achieve photorealism, although they struggle with high-resolution input images. To address this, Li et al. introduces PhotoWCT [38], which uses cascaded autoencoders to transform content features while keeping structural details intact. PhotoWCT2 [4] improves this by reducing parameters with blockwise training for faster stylization of high-resolution images. Wavelet-based WCT2 [39] uses wavelet pooling and un-pooling layers to preserve fine details during the STx process better.
Beyond feed-forward models, vision transformers excel at modeling long-range dependencies and spatial coherence, making them ideal for tasks where preserving local details and global structure is essential. StyTr2 by Deng et al. [7] is a prominent work that employs dual transformer encoders: one each for content and style; it introduces the idea of content-aware positional encoding (CAPE) to ensure scale invariance and spatial preservation. Contemporary works also explore unsupervised STx learning, with models like TransGAN [50], integrating transformers with GANs. More advanced attention mechanisms such as transformer-in-transformer (TNT) [48] further enhance STx by capturing complex dependencies through a nested architecture. Recent developments in diffusion-based models [32, 35] demonstrate further improvements by disentangling style and content for controlled (localized) adjustments.
On the other hand, physics-based methods use precise rendering to improve photorealism. For instance, Monte Carlo ray tracing simulates realistic lighting and shading to capture natural light interactions. Bidirectional reflectance distribution function (BRDF) modeling [57] preserves material properties like glossiness and roughness for more realistic appearance. These models consider how light reflects off surfaces to capture various light-material interactions. Another notable approach is deep image prior (DIP) [56], which uses a neural network to learn photorealism directly from the input image. It leverages natural image priors (  e.g., smooth regions, sharp edges) to handle tasks such as denoising, restoration, and style transfer. Moreover, retinex theory [91] powers many algorithms to improve photorealism by exploiting the geometry of how the human eye separates light from surface reflectance. Techniques like multi-scale retinex [92] restore true colors by ensuring natural color balance and contrast. Inverse tone mapping techniques [55] focus on rendering high-dynamic-range images, whereas optical flow-based STx [58] maintain spatial coherence to ensure smooth transitions in motion, shading, and texture gradients.
II-C Semantic and Multimodal Style Transfer
Semantic STx integrates a high-level understanding of image content to apply styles context-awarely, enhancing the coherence between the transferred style and the underlying content. For instance, Wang et al. [72] employs the SOTA semantic segmentation model DeepLab [71] to preserve object boundaries during STx. Lin et al. [68] further refine this approach by leveraging segmentation to maintain structural details, ensuring context-aware consistent STx. On the other hand, region-based models focus on selectively transferring styles to specific areas within an image. Niu et al. [74] propose a region-adaptive STx method that applies styles to distinct regions based on their semantic relevance. Liao et al. [77] develop a semantic-guided attention mechanism to enhance stylization in targeted regions, improving aesthetic quality and content preservation. These approaches allow for more precise control over the stylization process by fusing global semantic context and local features for improved STx.
Multimodal STx extends these concepts by incorporating additional modalities, such as depth information or textual descriptions, to guide the stylization. Ye et al. [70] explores multimodal learning techniques to fuse visual and textual data, enabling more expressive and versatile STx. Zheng et al. [69] applies retinal image segmentation combined with STx for medical imaging applications, demonstrating the versatility of multimodal approaches. Besides, few-shot and zero-shot learning techniques have been introduced to address the challenge of limited-style exemplars. Fu et al. [81] present StyleAdv, a few-shot adversarial learning framework that can adapt to new styles with minimal data. Huang et al. [82] proposes a zero-shot STx method using residual decay networks, eliminating the need for style images during training. Liu et al. [83] introduce StyleRF, which leverages neural radiance fields for zero-shot STx, enabling the application of novel styles without prior exposure.
Vision-language models (VLMs), particularly Contrastive Language–Image Pre-training (CLIP), enable the integration of textual and visual information for advanced image processing tasks [84]. Their robust image encoders capture comprehensive representations of image features, enhancing the potential for applications such as image STx. Clipstyler [9] framework achieves image STx solely from text descriptions by leveraging a patch-wise CLIP loss and perspective augmentation, thus eliminating the need for reference (style) images. Yang et al. [10] propose a zero-shot contrastive loss for diffusion models [33] that leverage patch-wise contrastive learning to preserve semantic content during text-guided STx without additional fine-tuning or auxiliary networks. These advancements significantly enhance the flexibility and scalability of STx models, making them more practical for real-world applications.
II-D Underwater Domain Style Transfer
Traditionally, domain STx methods are used to model underwater image distortions for synthetic data generation [93, 94]; the stylized distorted pairs are then used for learning image enhancement [95, 11] or improved object detection [15, 96]. These models enhance color correction, clarity, and stylization in a unified learning pipeline by separating content and style across different domains. For instance, Chen et al. [47] introduces a domain adaptation framework that leverages transformers for image enhancement [97, 98]. Zhou et al. [99] uses a transformer-based STx model for stylized representation learning of target recognition in underwater acoustic imagery. Our previous work AquaFuse [1] proposes a physics-guided method for geometry-preserving waterbody fusion for photometric data augmentation, which requires a prior reference image and known waterbody parameters. This work explores no-reference image STx from comprehensive multimodal learning databases – for the first time in the literature.
Figure 3: A few samples from the UF7D dataset are shown for each of its seven style categories.
III UF7D Dataset
III-A Data Preparation and Categorization
To enable data-driven learning of no-reference waterbody STx, we compile a comprehensive collection of high-resolution underwater image database named UF7D. It includes $7$ style categories: Blue (B), Blue-Green (B-G), Clear (C), Deep Blue (DB), Deep Green (DG), Green (G), and Green-Blue (G-B). A few samples for each category are shown in Figure 3.
(a) Distribution of images for all categories.
  Resolution # of Images
  $640\times 480$ $1633$
Up to $800\times 600$ $56$
Up to $1024\times 768$ $106$
Up to $1280\times 720$ $111$
Up to $1920\times 1080$ $374$
Above $1920\times 1080$ $1770$
Total # of Images 4050
(a) Image resolution information.
Figure 4: An overview of the UF7D dataset: it has a balanced distribution of seven waterbody style categories, each consisting of high-resolution natural underwater images.
A key feature of UF7D dataset is that it includes the scene depth maps for each image, generated by using the DepthAnything V2 model [100]. With high-resolution images and depth information, the UF7D dataset supports deep visual learning applications such as underwater image style transfer, enhancement, object detection, and more. To this end, our data collection process focused on the following aspects:
  1. $1$)
    Scene Diversity: UF7D includes a wide variety of natural underwater scenes that are sourced from various benchmark image databases [11, 12, 13, 101, 14, 15] and other public platforms such as Unsplash [102] and Flickr [103]. Our collection spans various objects including coral reefs, shipwrecks, pipelines, and various marine species (fish, plants, rocks) – enabling robust data-driven learning across diverse underwater conditions [104, 105].
  2. $2$)
    Image Quality: UF7D has a total of $4,050$ images, over $44\%$ of which are above $1920\times 1080$ resolution (see Fig. 4), offering high-resolution image contents necessary for fine-grained analyses.
  3. $3$)
    Balanced STx Categories: We relied on human visual perception for curating and compiling the seven style categories. As shown in Figure 4(a), these categories are evenly represented across the dataset, comprising the spectrum of dark blue to dark green colors, as well as clear images.
III-B Dataset Statistics and Analyses
Our statistical analysis reveals that UF7D consists of comprehensive coverage of varying underwater depths. Figure 5(a) illustrates the depth distributions by violin plots, where width represents the frequency density at each depth level. Here, we make the following observations.
  1. $\bullet$
    Cross-category Variation: Median depths range from $60$ to $116$ pixels across categories, with Green (G) environments showing the highest median and Deep Blue (DB) the lowest.
  2. $\bullet$
    Within-category Distribution: Categories exhibit distinct depth patterns; for instance, Clear (C) waters show consistent depth profiles, while Deep Blue (DB) shows a much wider variance.
(a) Violin plots showing depth value distributions across all categories. A particular example of depth variation is shown on the right; the blue box highlights a background region and the red box indicates a foreground region.
(b) PCA projection of global color characteristics (with $92\%$ variance captured in first two components).
(c) t-SNE visualization of local color statistics; the clusters validate the human-perceived category distinctions.
Figure 5: Statistical analyses and relevant information on the UF7D dataset.
We further perform detailed statistical analysis to validate our human-perceived style categorization process. Figure 5 presents the following two complementary analyses:
  1. $\bullet$
    Global Color Structure: Principle component analysis (PCA) reveals distinct clustering patterns for all categories, with the first two principal components capturing $92\%$ of total color variance; see Figure 5(b). The primary component embeds $76\%$ variance, effectively separating blue-dominated waters from green-dominated waters, which is consistent with our human perceptual evaluations.
  2. $\bullet$
    Local Color Relationships: A t-SNE (t-distributed stochastic neighbor embedding) visualization demonstrates clear separations among the style categories; see Figure 5(c). This confirms that images within each category share consistent spatial color characteristics.
These analyses validate that our waterbody style categorization based on human perception is statistically consistent as well. The inter-category and intra-category statistics also suggest that the underlying style can potentially be learned with tailored training strategies.
IV UStyle Model & Learning Pipeline
We present UStyle, a novel waterbody style transfer framework that integrates a deep encoder-decoder architecture, progressive blockwise training, and a depth-aware stylization module. Given a content image $\mathbf{I}_{c}\in\mathbb{R}^{H\times W\times 3}$ and a style image $\mathbf{I}_{s}\in\mathbb{R}^{H\times W\times 3}$, UStyle generates a stylized output $\mathbf{I}_{cs}$ by fusing waterbody characteristics from the style image while preserving scene geometry of the content image. The depth maps $\mathbf{D}_{s}$ and $\mathbf{D}_{c}$, representing the depth information of the style and content images, respectively, adapt features based on depth. $\mathbf{D}_{s}$ guides the extraction of waterbody details from the style image, while $\mathbf{D}_{c}$ ensures content-aware stylization by preserving structural consistency and depth-dependent variations during style transfer. As shown in Fig. 6, the UStyle learning pipeline comprises three major components: (1) an encoder-decoder backbone for multi-scale feature extraction with fine spatial details and high-level semantics; (2) a progressive training strategy for stable coarse-to-fine stylization; and (3) a depth-aware stylization module for effective waterbody style transfer.
Figure 6: Overview of our UStyle model, featuring an encoder-decoder backbone with skip connections, high-frequency residual (HFR) paths, and a CLIP-based coarse similarity module for aligning intermediate features with semantic cues. A multi-stage learning strategy is adopted that progressively trains deeper symmetric blocks to preserve both coarse structure and fine details in the waterbody fusion process.
IV-A Encoder–Decoder Backbone
We use a ResNet50 [18] encoder ($\mathcal{E}$) partitioned into five sequential blocks $\mathrm{EB}_{1:5}$, which extract increasingly finer features $\{{F}_{i,c}\}_{i=0:4}$ and $\{{F}_{i,s}\}_{i=0:4}$ from a given content and style images, respectively. These feature maps are then fed into a decoder $\mathcal{D}$ with hierarchical ($\mathrm{EB}_{i-1}\to\mathrm{DB}_{i}$) skip connections. The symmetric skip connections retain spatial structures and propagate high-frequency details from the encoder to the decoder. Each decoder block $\mathrm{DB}_{i}$ consists of a transposed convolution [106] for upsampling, followed by a convolutional layer [107], batch normalization [108], and a ReLU activation [109]. Rather than explicitly computing residuals via average pooling and subtraction, our design preserves these details through the original skip connections. These features are fused to achieve waterbody STx with pixel-wise guidance from the content depth map $\mathbf{D}_{c}$, ensuring that the fused features ${F}_{i,cs}$ adapt to both structural (object geometry) and stylistic (waterbody) elements. This fusion is performed by our novel depth-aware stylization module; see Section IV-C.
IV-B Progressive Pre-Training and Domain-specific Finetuning
We implement a sequential blockwise approach in UStyle, where the encoder-decoder network is optimized progressively over five stages ($N=1$:$5$). Completed $\mathrm{EB}_{i}-\mathrm{DB}_{i}$ blocks are frozen to ensure stability of the training, as shown in Fig. 6. Initially, the model is pre-trained on a large-scale atmospheric dataset (MSCOCO-2017 [110]) with a weighted combination of reconstruction loss [111], feature reconstruction loss [86], and perceptual loss. During the training of the first stage ($N=0$), the network is optimized using only a reconstruction loss $\mathcal{L}_{r}=\|\mathbf{I}_{c}-\hat{\mathbf{I}}_{c}\|_{2}^{2}$ where $\hat{\mathbf{I}}_{c}$ denotes the decoder’s output. In subsequent stages ($N>0$), additional losses are introduced to ensure feature-level consistency and perceptual fidelity. These include a feature reconstruction loss $\mathcal{L}_{feat}=\|\mathbb{F}_{i}-\hat{\mathbb{F}}_{i}\|_{2}^{2}$ and a perceptual loss $\mathcal{L}_{percept}=\|\phi(\mathbf{I}_{c})-\phi(\hat{\mathbf{I}}_{c})\|_{2}^{2}$, where $\phi(\cdot)$ is a high-level feature extractor based on VGG-16 [24].
After pretraining, UStyle is fine-tuned on a large-scale underwater dataset (USOD10K [112]) for domain-specific supervision with SSIM loss $\mathcal{L}_{ssim}=1-SSIM(\mathbf{I}_{c},\hat{\mathbf{I}}_{c})$, color consistency loss in LAB image space $\mathcal{L}_{color}=\|\mathrm{LAB}(\mathbf{I}_{c})-\mathrm{LAB}(\hat{\mathbf{I}}_{c})\|_{2}^{2}$, and frequency-domain loss by Fast Fourier Transforms (FFT) $\mathcal{L}_{fft}=\|\mathcal{F}(\mathbf{I}_{c})-\mathcal{F}(\hat{\mathbf{I}}_{c})\|_{2}^{2}$. Moreover, CLIP-based content supervision is incorporated by introducing a CLIP loss [25] to ensure semantic alignment. It is defined as:
$\mathcal{L}_{clip}=1-\mathrm{cosine\_similarity}\Bigl{(}\mathrm{CLIP}({\mathbf{I}}_{c}),\mathrm{CLIP}(\hat{\mathbf{I}}_{c})\Bigr{)}.$ (1)
The overall objective functions for multi-stage training are as follows:
$\mathcal{L}_{0}$ $=\mathcal{L}_{r}+\mathcal{L}_{ssim}+\mathcal{L}_{color}+\mathcal{L}_{fft}+\mathcal{L}_{clip}\quad(N=0)$ (2)
$\mathcal{L}_{N}$ $=\mathcal{L}_{r}+\mathcal{L}_{ssim}+\mathcal{L}_{color}+\mathcal{L}_{fft}+\mathcal{L}_{clip}+\mathcal{L}_{feat}+\mathcal{L}_{percept}\quad(N=1,2,3,4)$ (3)
This progressive training approach allows the network to learn robust multi-scale feature representations while ensuring content, style, and depth consistency for high-fidelity underwater waterbody style transfer.
IV-C Depth-aware Stylization
We design a depth-aware stylization function: DA-WCT which extends the multi-scale blending concept from the WCT (whitening and coloring transform) literature [37]. DA-WCT integrates a physics-guided waterbody estimation method for depth-guided feature blending via WCT. First, we estimate the waterbody from the style image $\mathbf{I}_{s}$ using a physics-guided method [1]; from its depth map $\mathbf{D}_{s}$, we leverage the $5\%$ farthest pixels’ median values to estimate the waterbody background in $\mathbf{I}_{s}$. A color filter is then applied to retain only bluish-green pixels, removing red-channel influence due to differential light attenuation underwater. The extracted waterbody $\mathbf{B}_{s}$ is then fed into the encoder to obtain the style features.
Let $F_{c}$ and $F_{s}$ denote the content and style feature maps extracted from the encoder, respectively. We first perform a WCT-based stylization to align the statistics of $F_{c}$ with those of $F_{s}$; the stylized feature $\tilde{F}_{cs}$ is computed as $\tilde{F}_{cs}\;=\;\mu_{s}\;+\;C_{s}^{1/2}\,C_{c}^{-1/2}\,\small(F_{c}-\mu_{c}\small)$, with $\mu_{c}$, $\mu_{s}$ being the channel-wise means and $C_{c}$, $C_{s}$ the covariance matrices of $F_{c}$ and $F_{s}$, respectively. Let $\mathbf{D}_{c}$ denote the single-channel content depth map, a global depth threshold $\tau\in[0,1]$ is computed using Otsu’s method [113]. Next, an adaptive scaling factor $k$ is derived by first smoothing $\mathbf{D}_{c}$ via average pooling with a $5\times 5$ kernel, then computing the global mean and standard deviation as follows:
$\mu_{d}$ $=\frac{1}{H\cdot W}\sum_{x}\mathrm{avgpool}\big{(}\mathbf{D}_{c}(x)\big{)}$ (4)
$\sigma_{d}$ $=\sqrt{\frac{1}{H\cdot W}\sum_{x}\Bigl{(}\mathrm{avgpool}\big{(}\mathbf{D}_{c}(x)\big{)}-\mu_{d}\Bigr{)}^{2}}$ (5)
With the tunable scaling factors $\tau$ and $k$ determined, the depth weight map for each pixel is computed using the sigmoid function.
$w(x)=\sigma\Bigl{(}-k\cdot\bigl{(}\mathbf{D}_{c}(x)-\tau\bigr{)}\Bigr{)}=\frac{1}{1+\exp\Bigl{(}k\cdot\bigl{(}\mathbf{D}_{c}(x)-\tau\bigr{)}\Bigr{)}}$ (6)
Subsequently, the blended feature is obtained by merging the stylized features $\tilde{F}_{cs}(x)$ with the original content features $F_{c}(x)$ as follows
$F_{\text{depth}}(x)=w(x)\cdot\tilde{F}_{cs}(x)+[1-w(x)]\cdot F_{c}(x)$ (7)
In our formulation, pixels with higher depth values (i.e., farther objects) yield $w(x)\rightarrow 1$ and favor the stylized features $\tilde{F}_{cs}(x)$. In contrast, pixels with lower depth values get $w(x)\rightarrow 0$, preserving more of the original content features $F_{c}(x)$. A global style strength parameter $\alpha\in[0,1]$ is applied to control the degree of the underlying stylization process.
$F_{cs}\;=\;\alpha\,F_{\text{depth}}\;+\;\bigl{(}1-\alpha\bigr{)}\,F_{c}.$ (8)
[…]
$\mathbf{I}_{stylized}=\text{GIF}(\mathbf{I}_{cs})=\text{GIF}\big{(}\mathcal{D}(F_{\text{multi}})\big{)}.$ (9)
Figure 7: Inference pipeline of the proposed UStyle model. The content and style images and their corresponding depth maps are first processed through a ResNet-based encoder to extract multi-scale features. These features are processed by our depth-aware whitening and coloring transform (DA-WCT) function for waterbody style alignment, which are then progressively decoded using a multi-stage decoder. Finally, the stylized output is refined with a guided filter to improve visual consistency.
Implementation Details. UStyle is implemented in PyTorch with OpenCV and Torch libraries. We use Adam optimizer with an initial learning rate of $10^{-4}$ for the supervised learning. Training is conducted on a Linux system equipped with $4$ GPUs and $32$ GB of RAM. At the first stage of training, the encoder-decoder network is pre-trained on the MSCOCO-2017 with $118$ K training samples for $10$ epochs. It is then fine-tuned on USOD10K underwater dataset with $10$ K examples for another $10$ epochs. The whole training and fine-tuning takes $82$ hours to complete. Once trained, the frozen graph is used for inference, executed on a single GPU and $16$ GB of RAM.
Inference Pipeline. The inference process of UStyle follows a structured pipeline to achieve high-fidelity underwater style transfer while maintaining the structural integrity of the content image. As illustrated in Fig. 7, the process begins with a given content image $\mathbf{I}_{c}$ and a style image $\mathbf{I}_{s}$, along with their corresponding depth maps $\mathbf{D}_{c}$ and $\mathbf{D}_{s}$. The style image is first processed by our physics-guided Waterbody Extractor [1], which isolates the dominant waterbody characteristics $\mathbf{B}_{s}$. Both $\mathbf{I}_{c}$ and $\mathbf{B}_{s}$ are then passed into the Encoder, which extracts multi-scale feature representations $\mathbf{F}_{c}$ and $\mathbf{F}_{s}$, respectively. The extracted features are then processed by our DA-WCT module, which aligns the feature statistics between content and style domains, producing stylized feature representations $\mathbf{F}_{cs}$. These transformed features are fed into a multi-stage Decoder, each progressively refining the stylized output. Skip connections from the Encoder to the Decoder help retain spatial consistency and fine-grained structural details. At each decoding stage, an adaptive fusion parameter $\alpha$ controls the style influence level, ensuring smooth content-style integration. Finally, the stylized image $\mathbf{I}_{cs}$ is passed through a Guided Filter, which enhances structural coherence and removes unwanted artifacts, yielding the final stylized output $\mathbf{I}_{stylized}$.
V Experimental Results
V-A Baseline Models and Metrics
We use six SOTA models from artistic and photorealistic STx domains as baselines for performance evaluation. Our comparison includes artistic STx methods StyTR-2 [115], ASTMAN [17], MCCNet [16], and IEContraAST [116], alongside the photorealistic method PhotoWCT2 [4]. Additionally, we include AquaFuse [1], a physics-based underwater waterbody style transfer method, to benchmark our approach further. For a fair comparison, all models are fine-tuned on the UF7D dataset; specifically, their pre-trained weights are fine-tuned over $160$ K iterations with a batch size of $8$ while keeping all other hyperparameters at their recommended settings. For evaluation, we define four cross-domain underwater style transfer test cases: Blue (B) $\rightarrow$ Green (G), Green (G) $\rightarrow$ Blue (B), Deep Blue (DB) $\rightarrow$ Deep Green (DG), and Deep Green (DG) $\rightarrow$ Deep Blue (DB). Each case consists of $25$ content images and $25$ style images, generating $25\times 25=625$ unique test scenes per case, resulting in a total of $4\times 625=2,500$ test scenes.
Evaluation Metrics. We use five standard metrics to evaluate the performance of UStyle and other baseline methods. These complementary metrics assess both the structural preservation and perceptual quality of the stylized images generated by each model. The metrics are:
  • Root Mean Square Error (RMSE) [117] quantifies pixel-level accuracy between stylized and content images as: ${\tt{RMSE}}=\sqrt{\tt{MSE}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(x_{i}-y_{i})^{2}}$; where $x_{i}$ and $y_{i}$ are corresponding pixel values and $N$ is total number of pixels. The squared difference term amplifies larger errors, making RMSE especially sensitive to outliers and structural distortions.
  • Peak Signal-to-Noise Ratio (PSNR) [22] measures reconstruction quality as ${\tt{PSNR}}=10\cdot\log_{10}\left(\frac{L^{2}}{\tt{MSE}}\right)$; where $L=255$ represents the maximum possible pixel value in 8-bit images, and MSE is the mean squared error. PSNR’s logarithmic scale and dependency on $L$ make it particularly sensitive to illumination changes and global color shifts.
  • Structural Similarity Index (SSIM) [118] evaluates the geometric consistency of the generated image based on luminance, contrast, and changes in structural information. It is calculated as:
    ${\tt{SSIM}}(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})}$ (10)
    where $\mu$ and $\sigma$ represent mean and variance respectively, and $C_{1}$, $C_{2}$ are stability constants. The multiplicative combination of luminance ($\mu$), contrast ($\sigma$), and structure ($\sigma_{xy}$) terms makes SSIM particularly effective for evaluating perceptual image quality.
  • Gradient Magnitude Similarity Deviation (GMSD) [119] assesses edge preservation by exploiting image gradient information as: ${\tt{GMSD}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(g_{i}-\mu_{g})^{2}}$; where $g_{i}$ represents gradient magnitude similarity at pixel $i$, and $\mu_{g}$ is the mean. GMSD is particularly effective at detecting structural and textural variations while being relatively insensitive to global illumination changes.
  • Learned Perceptual Image Patch Similarity (LPIPS) [120] measures perceptual similarity as:
    ${\tt{LPIPS}}(x,y)=\sum_{l}\frac{1}{H_{l}\,W_{l}}\sum_{h,w}\Big{\|}w_{l}\odot(\phi_{l}(x)_{h,w}-\phi_{l}(y)_{h,w})\Big{\|}_{2}^{2}$ (11)
    Here, $\phi_{l}$ represents feature maps from layer $l$ of a pre-trained network, and $w_{l}$ are learned weights that calibrate the importance of different network layers. This metric captures high-level perceptual similarities that often correlate better with human perceptual judgment than traditional metrics.
This combination of metrics enables the evaluation of both low-level image statistics and high-level perceptual qualities. All evaluations are conducted on images normalized to $640\times 480$ resolution using their official metric implementations.
V-B Qualitative Results
[…]
Figure 8: Qualitative comparison of underwater image style transfer (STx) results across different methods. The first two columns show content and style images, followed by stylized outputs from StyTR-2 [7], ASTMAN [17], MCCNet [6], IEContraAST [5], PhotoWCT2 [4], AquaFuse [1], and UStyle. Each row shows two examples of: B $\rightarrow$ G, G $\rightarrow$ B, DB $\rightarrow$ DG, and DG $\rightarrow$ DB transformations.
V-C Quantitative Results
We conduct a comprehensive quantitative evaluation using five complementary metrics across various underwater STx scenarios from the UF7D test set. As presented in Table I(a), UStyle achieves 4×4$\times$4 × to 6×6$\times$6 × lower RMSE rates across all scenarios compared to SOTA artistic and photorealistic STx methods. It also consistently outperforms competing approaches in PSNR and SSIM, as shown in Table I(b). The closest competitor, AquaFuse, requires a carefully selected reference image and prior information such as incident angle and attenuation characteristics. In contrast, UStyle achieves improvements of $6.8\%$ (RMSE), $2.1\%$ (PSNR), and $3.3\%$ (SSIM) over AquaFuse while operating without any reference or prior knowledge — demonstrating its robustness and accuracy for waterbody STx.
[…]
TABLE I: Quantitative comparison of averaged RMSE ($\downarrow$), PSNR ($\uparrow$), and SSIM ($\uparrow$) across different underwater STx methods are conducted on $2,500$ test samples in UF7D. The best score is highlighted in bold blue, while the second-best is shown in blue.
(a) RMSE $\downarrow$ B $\rightarrow$ G G $\rightarrow$ B DB $\rightarrow$ DG DG $\rightarrow$ DB
ASTMAN [17] $52.29$ $54.53$ $48.91$ $63.74$
IEContraAST [116] $47.32$ $50.54$ $58.17$ $69.49$
MCCNet [16] $44.64$ $53.12$ $46.04$ $60.71$
PhotoWCT2 [4] $34.68$ $34.80$ $31.89$ $26.66$
StyTR-2 [115] $37.08$ $40.61$ $47.18$ $54.01$
AquaFuse [1] $10.45$ $10.19$ $10.15$ $10.24$
UStyle $9.39$ $9.35$ $9.50$ $8.96$
(b) PSNR $\uparrow$ B $\rightarrow$ G G $\rightarrow$ B DB $\rightarrow$ DG DG $\rightarrow$ DB
ASTMAN [17] $14.95$ $14.37$ $14.99$ $13.16$
IEContraAST [116] $15.90$ $14.40$ $13.82$ $12.10$
MCCNet [16] $15.81$ $14.01$ $15.63$ $13.35$
PhotoWCT2 [4] $19.75$ $18.70$ $18.95$ $20.05$
StyTR-2 [115] $17.86$ $16.67$ $15.59$ $14.58$
AquaFuse [1] $27.77$ $28.00$ $28.03$ $27.94$
UStyle $28.73$ $28.72$ $28.64$ $29.14$
(c) SSIM $\uparrow$ B $\rightarrow$ G G $\rightarrow$ B DB $\rightarrow$ DG DG $\rightarrow$ DB
ASTMAN [17] $0.8026$ $0.6334$ $0.8150$ $0.6397$
IEContraAST [116] $0.7641$ $0.6304$ $0.7887$ $0.6033$
MCCNet [16] $0.8095$ $0.6302$ $0.8290$ $0.6570$
PhotoWCT2 [4] $0.8675$ $0.7819$ $0.8324$ $0.8030$
StyTR-2 [115] $0.8494$ $0.6937$ $0.8199$ $0.7125$
AquaFuse [1] $0.9337$ $0.8708$ $0.7394$ $0.8812$
UStyle $0.9613$ $0.8818$ $0.8389$ $0.9126$
(b) (d) Average RMSE scores are shown on (top barchart) and the PSNR--SSIM scores are visualized in a 2D grid (bottom); evident from these results, UStyle achieves better overall performance than existing STx methods.
TABLE II: Quantitative comparison of averaged GMSD ($\downarrow$) and LPIPS ($\downarrow$) across different underwater STx methods, conducted on $2,500$ test samples in UF7D. The best score is highlighted in bold blue, while the second-best is shown in blue.
(a) GMSD $\downarrow$ B $\rightarrow$ G G $\rightarrow$ B DB $\rightarrow$ DG DG $\rightarrow$ DB
ASTMAN [17] $0.0693$ $0.0318$ $0.0508$ $0.0179$
IEContraAST [116] $0.0694$ $0.0300$ $0.0556$ $0.0187$
MCCNet [16] $0.0262$ $0.0107$ $0.0386$ $0.0386$
PhotoWCT2 [4] $0.0622$ $0.0725$ $0.0762$ $0.0386$
StyTR-2 [115] $0.0551$ $0.0287$ $0.0380$ $0.0187$
AquaFuse [1] $0.0252$ $0.0631$ $0.0201$ $0.0471$
UStyle $0.0227$ $0.0564$ $0.0200$ $0.0431$
(b) LPIPS $\downarrow$ B $\rightarrow$ G G $\rightarrow$ B DB $\rightarrow$ DG DG $\rightarrow$ DB
ASTMAN [17] $0.2992$ $0.4849$ $0.3769$ $0.4091$
IEContraAST [116] $0.2762$ $0.4835$ $0.3896$ $0.4518$
MCCNet [16] $0.2838$ $0.4976$ $0.3677$ $0.3989$
PhotoWCT2 [4] $0.1676$ $0.2618$ $0.2409$ $0.1755$
StyTR-2 [115] $0.2372$ $0.3612$ $0.3624$ $0.3261$
AquaFuse [1] $0.4274$ $0.3902$ $0.5597$ $0.3789$
UStyle $0.1926$ $0.2235$ $0.1342$ $0.1373$
(a) (c) A heatmap illustrating the combined deviation, which represents the overall distortion by each model, computed by GMSD and LPIPS scores. Here, lower values indicate better overall performance; UStyle achieves the lowest score, $19.8\%$ lower than the closest competitor.
V-D Ablation Studies
We conduct ablation experiments to evaluate the key components of UStyle: (i) the impact of depth supervision in the stylization process and (ii) the contribution of various loss components to STx performance. To assess the role of depth supervision, we compare stylized results with and without depth guidance. As shown in Fig. 9, DA-WCT improves structural retention and color adaptation in complex underwater scenes. Without depth supervision, stylized images exhibit color distortions and detail loss, whereas DA-WCT preserves content integrity throughout the STx process.
Additionally, we investigate the contribution of different loss components by evaluating three configurations of our model: (1) UStyle without clip loss (w/o CLIP), (2) UStyle without additional regularization losses, $L3=\mathcal{L}_{ssim}+\mathcal{L}_{color}+\mathcal{L}_{fft}$ (w/o $L3$), and (3) UStyle with all loss components. Table III presents the quantitative comparison across PSNR, SSIM, and RMSE for four underwater style transfer scenarios. The results show that the full UStyle model consistently performs better, demonstrating the effectiveness of depth supervision and loss regularization. From a quantitative perspective, depth supervision leads to a substantial improvement in all three metrics. The inclusion of additional losses further refines the model’s ability to balance structural preservation and perceptual realism. The full UStyle model achieves the highest PSNR and SSIM values while maintaining the lowest RMSE across all scenarios.
To further examine the impact of DA-WCT stylization, we employ Grad-CAM [121] visualizations to highlight the regions of interest in our model’s decision-making process. As illustrated in Fig. 10, depth supervision enables UStyle to focus more effectively on salient structures, ensuring spatially coherent style adaptation. The middle row displays activation maps generated with depth guidance, where attention is evenly distributed across foreground and background elements, enhancing structure retention. In contrast, the bottom row—corresponding to the model without depth supervision—reveals inconsistent attention patterns, leading to fine-detail loss and exaggerated distortions. Notably, DA-WCT ensures that complex structures, such as marine organisms and underwater artifacts, are better preserved by aligning style fusion with content depth information. These visualizations further validate the importance of incorporating depth guidance, reinforcing its role in improving geometric consistency and reducing artifacts in waterbody STx.
These findings confirm that depth supervision is essential for maintaining spatial coherence and content preservation in underwater STx. Furthermore, integrating additional regularization losses significantly enhances perceptual quality, reducing distortions and improving color adaptation.
Figure 9: A few ablation results illustrating the impact of depth-guidance in the stylization process. With the proposed DA-WCT blending, stylized outputs exhibit consistent blending, both on foreground objects and background waterbody regions. In contrast, we observe a global averaging effect while using only WCT and over-saturated pixels for deep-water (DB/DG) styles.
TABLE III: For ablation study, we compare the performance of three variants: UStyle without clip loss (w/o CLIP), UStyle without additional regularization losses $L3$ (w/o $L3$), and UStyle with all loss components. Each of these variants is evaluated with WCT/DA-WCT on PSNR (↑), SSIM (↑), and RMSE (↓) metrics.
Data B $\rightarrow$ G G $\rightarrow$ B DB $\rightarrow$ G DG $\rightarrow$ B
Metric Model WCT DA-WCT WCT DA-WCT WCT DA-WCT WCT DA-WCT
PSNR 1: UStyle (w/o CLIP) $28.03$ $28.69$ $28.24$ $28.69$ $27.85$ $28.00$ $28.39$ $29.09$
2: UStyle (w/o $L3$) $27.96$ $28.36$ $28.21$ $28.64$ $27.84$ $27.91$ $28.43$ $28.97$
3: UStyle (Full) $28.09$ $28.73$ $28.24$ $28.72$ $27.91$ $28.64$ $28.43$ $29.14$
SSIM 1: UStyle (w/o CLIP) $0.9319$ $0.9549$ $0.8090$ $0.8758$ $0.6365$ $0.8046$ $0.8421$ $0.9102$
2: UStyle (w/o $L3$) $0.9337$ $0.9590$ $0.8004$ $0.8675$ $0.6087$ $0.7780$ $0.8388$ $0.9069$
3: UStyle (Full) $0.9349$ $0.9613$ $0.8260$ $0.8818$ $0.6727$ $0.8389$ $0.8506$ $0.9126$
RMSE 1: UStyle (w/o CLIP) $10.15$ $9.43$ $9.90$ $9.40$ $10.37$ $10.19$ $9.75$ $9.01$
2: UStyle (w/o $L3$) $10.23$ $9.77$ $9.93$ $9.45$ $10.38$ $10.28$ $9.70$ $9.15$
3: UStyle (Full) $10.08$ $9.39$ $9.89$ $9.35$ $10.31$ $9.50$ $9.70$ $8.96$
Figure 10: Grad-CAM heatmap visualizing the effect of depth-aware style fusion in UStyle. Our DA-WCT enhances attention to structural details, ensuring a more spatially coherent style adaptation (third row). In comparison, without depth supervision, attention is unevenly distributed, leading to potential distortions and loss of fine details in complex underwater scenes (second row).
VI Analysis and Discussion
VI-A Analysis of Results
[…]
Complementary to these numerical results, the Grad-CAM [121] visualizations in Figure 10 illustrate that depth guidance enables more uniform and focused attention across critical image regions, ensuring spatial coherence. These findings suggest that the DA-WCT module guides the network to focus on relevant salient structural cues while mitigating artifacts. Moreover, the training process revealed that UStyle consistently converges across diverse underwater conditions without manual parameter tuning, unlike physics-based models, which are sensitive to reference image quality and scene priors. Overall, UStyle’s design choices yield a robust and efficient framework for underwater STx, and these insights may guide future research in developing even more adaptive and efficient solutions.
Furthermore, during inference using UStyle and the baseline models, we noticed that models without depth guidance, mainly when applied to deep-water style transitions, frequently produced over-saturated outputs with a global averaging effect [122]; this resulted in increased RMSE values by approximately $10\%$ to $12\%$ and LPIPS scores that were consistently higher. In contrast, UStyle maintained stable convergence and demonstrated a more consistent loss decay pattern with a reduction in loss variance of about $15\%$. Furthermore, the inclusion of the CLIP loss [25] in UStyle contributed to a smoother convergence trajectory by aligning semantic features; removal of this component led to slight degradations in SSIM (a reduction of roughly $0.002$ on average) and a marginal increase in RMSE, highlighting the importance of semantic consistency in waterbody STx learning.
Additional observations from our experiments reveal that UStyle’s adaptive multi-stage training strategy [20] resulted in convergence speeds approximately $20\%$ faster than several baseline models. The fine-tuning process on underwater-specific data proved crucial for the network to adapt to domain-specific distortions, yielding significant improvements in performance metrics across diverse underwater scenarios. This indicates that carefully designed loss functions, particularly those focusing on structural similarity [22], color consistency [21], and frequency-domain fidelity [23], are essential for achieving high-quality results in complex underwater environments.
During waterbody fusion, we observed that the depth maps’ quality and depth thresholding precision influenced the stylization process. Variations in depth maps lead to noticeable differences in the retention of fine details, especially in regions with subtle gradient changes. Our future research will emphasize robust depth extraction methods and consider dynamic adjustment of depth-based parameters to enhance performance further. Additionally, the success of UStyle in handling diverse underwater conditions without reliance on manually curated reference images with explicit scene priors suggests that its framework may be extended to other domains where scene priors are challenging to obtain.
VI-B Computational Analyses
As shown in Table IV, UStyle maintains a stable inference rate across multiple resolutions. Evaluated on a single NVIDIA A100 GPU, UStyle processes a $256\times 256$ image in approximately $518$ ms and a full HD $1920\times 1080$ image in about $741$ ms. While transformer-based methods such as StyTr2 [7] and IEContraAST [5] struggle with high-resolution images due to high memory consumption, UStyle consistently operates across all resolutions, demonstrating scalability. Although MCCNet [115] and ASTMAN [17] achieve faster processing times at low resolutions, their performance is not comparable to UStyle. Moreover, due to memory constraints, StyTr2 [7], ASTMAN [17], and IEContraAST [5] fail to process images at resolutions beyond $1280\times 720$ as well. While UStyle is not intended for real-time applications, rather a high-performance data augmentation and stylazation tool, it still provides competitive processing speeds, especially at high input resolutions.
TABLE IV: Inference speed based on image processing time per image (msec $\downarrow$) across different image resolutions for six data-driven SOTA methods are shown. Evaluations are performed on a single NVidia A100 GPU with $16$ GB RAM; ‘$\times$’ indicates the out-of-memory (OOM) cases.
Resolution $256\times 256$ $512\times 512$ $640\times 480$ $1024\times 1024$ $1280\times 720$ $1920\times 1080$ $3840\times 2160$
ASTMAN [17] $180$ $190.8$ $222.7$ $250.6$ $227.8$ $274.0$$\times$
IEContraAST [5] $1219.5$ $2127.7$ $2272.7$ $2702.7$ $3125.0$ $4761.9$$\times$
MCCNet [6] $170.9$ $209.2$ $229.4$ $354.6$ $371.7$ $549.5$ $1886.8$
StyTr2 [7] $136.2$ $274.7$ $285.7$ $1639.3$ $2040.8$$\times$ $\times$
PhotoWCT2 [4] $323.6$ $340.1$ $344.8$ $390.6$ $374.5$ $467.3$ $885.0$
UStyle $518.1$ $621.1$ $641.0$ $645.2$ $645.2$ $740.7$ $1087.0$
VI-C Benefits Over AquaFuse
As we discussed earlier in Sec. IV-C, we use a physics-guided waterbody synthesis module of AquaFuse [1] to extract the waterbody characteristics from the style image. Note that AquaFuse synthesizes the content image by incorporating a reference scene’s waterbody through a physics-guided approximation approach. However, this method relies on a noise-free reference image and prior scene information, including depth, incident angle, and attenuation parameters. As illustrated in Figure 11, AquaFuse is highly sensitive to input parameters such as incident angles and tuned attenuation coefficients, limiting its generalizability across diverse underwater conditions. Its performance is strongly influenced by the choice of reference image and variations in incident angles. In contrast, UStyle, a learning-based STx framework, adapts robustly to varying aquatic environments without requiring manually tuned priors. To facilitate this adaptability, large-scale datasets such as USOD10K and UF7D provide a diverse collection of waterbody styles, enabling more generalized STx learning, as demonstrated in our experimental analyses.
Figure 11: Comparison between AquaFuse and UStyle across different underwater scenes. The first three rows show that AquaFuse outputs are highly sensitive to the choice of incident angle; the last row further shows that selecting a wrong reference image can significantly impact its waterbody fusion process. UStyle is invariant to these dependencies, offering consistent waterbody STx performance.
VI-D Limitations and Future Directions
Although UStyle demonstrates robust performance and achieves SOTA results, several aspects of the framework warrant further refinement. In particular, enhancements in depth estimation accuracy, training efficiency, adaptive handling of specific waterbody styles, and performance under extreme turbidity conditions remain as challenging cases; more details are listed below.
  1. $\bullet$
    Dependency on depth estimation: UStyle’s performance, as dependent on the DA-WCT module, depends on the quality of depth maps, which can be unreliable in noisy sensing conditions and severely attenuated scenes. Future work could incorporate self-supervised depth estimation methods to enhance robustness and reduce dependency on precomputed scene depth estimation.
  2. $\bullet$
    Training and inference efficiency: Although the progressive blockwise training strategy enhances convergence stability and multi-scale feature learning, it increases training time compared to single-stage models. Future work could explore the design of lightweight architectures or leverage knowledge distillation techniques [123] to improve efficiency without sacrificing the generalization performance.
  3. $\bullet$
    Extremely turbid or rare waterbody styles: The compounded effects of scattering and absorption can significantly impair style transfer quality for severely turbid scenes [124]. Future work will explore integrating polarization-based enhancement techniques [125] and multi-spectral image enhancement methods [126] to improve UStyle’s robustness in challenging underwater environments. We will also explore adaptive blending mechanisms [127] to better accommodate rare/unseen waterbody styles.
Despite these challenges, UStyle significantly advances underwater STx literature for data-driven waterbody fusion. Future research will optimize its efficiency, extend it to real-time video applications, and improve robustness in complex underwater scenarios.
VII Conclusion
In this work, we presented UStyle, the first data-driven framework for underwater waterbody style transfer that robustly fuses waterbody characteristics without relying on manually curated reference images with explicit scene priors. Leveraging a ResNet-based encoder-decoder architecture, a progressive blockwise training strategy, and a novel depth-aware whitening and coloring transform (DA-WCT), UStyle effectively preserves structural integrity while achieving perceptually compelling stylization. Extensive quantitative and qualitative evaluations demonstrate that UStyle outperforms SOTA artistic, photorealistic, and physics-based methods, achieving improvements of up to $16\%$ in RMSE, reaching PSNR values of $29.14$  dB, SSIM values as high as $0.9613$, and attaining LPIPS scores as low as $0.1373$. Ablation studies further confirm that integrating depth supervision and additional regularization losses is critical for enhancing geometric consistency and minimizing perceptual distortions. Moreover, computational analyses reveal that UStyle maintains stable inference rates across various resolutions, showcasing its scalability for high-resolution underwater imaging applications. Although challenges remain in handling extreme turbidity and further optimizing training efficiency, our findings underscore the potential of UStyle as a robust tool for underwater imaging, data augmentation, and vision tasks. Future work will incorporate depth estimation, explore lightweight architectures for real-time applications, and extend the framework to handle more complex underwater STx scenarios in real-time.
Acknowledgments
This work is supported in part by the National Science Foundation (NSF) grant #$2330416$ and the University of Florida (UF) research grant #$132763$.